OpenStudy (anonymous):

A 60-kg man standing on a scale in an elevator notes that as the elevator rises, the scale reads 819 N. What is the acceleration of the elevator?

OpenStudy (anonymous):

\[\Sigma F_{up}=ma\\ N-mg=ma\\a=\frac{N-mg}{m} \] Whwre N is the scale reading.

OpenStudy (anonymous):

he is correct n = 819

OpenStudy (anonymous):

so then mg is 60 how do i get m?

OpenStudy (anonymous):

mg=60*9.8

OpenStudy (anonymous):

thanks i just thought of that lol